OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues.

OpenStudy (anonymous):

-9 - 4 + 1 + 6 + ... + 66

OpenStudy (anonymous):

you can use the formula for sums of geometric and arithmetic series.

OpenStudy (anonymous):

Sum of arithmetic finite series: \[S=(a_1+a_k)*k/2\]

OpenStudy (anonymous):

what does k mean in the formula

OpenStudy (anonymous):

k means the number of last term, which is in your sequence is 66. u have an arithmetic series that starts with term -9 and has difference 5. Here is your sequence: \[a_n=-9+5(n-1)\]

OpenStudy (anonymous):

so k will mean here, the number of term which is equal to 66. To find k do the following: \[66=-9+5(k-1)\] \[66+9=5(k-1)\] \[75/5=k-1\] \[15+1=k=16\]

OpenStudy (anonymous):

so the last term is 6th and you can substitute this number into formula \[ S=(-9+66)*6/2=57*3=171\]

OpenStudy (anonymous):

the answer is 171, @allydiaz

OpenStudy (anonymous):

All the options are in summation notation @ksanka

OpenStudy (anonymous):

How would I write it in summation notation form?

jimthompson5910 (jim_thompson5910):

nth term a1 + d(n-1) -9 + 5(n-1) -9 + 5n - 5 5n - 14 so the general nth term is 5n - 14 we'll use k instead of n since that's common for summations. So we'll have 5k - 14 instead The summation will look like this \[\LARGE \sum_{k=1}^{16}(5k-14)\] the funky looking E, written like this \(\Large \Sigma\) is the capital Greek letter sigma. Sigma for Sum the "k = 1" portion tells us where to start (start at the first term) the 16 up top tells us where to stop: we stop when we get to the 16th term the (5k-14) off to the right tells us what we're actually adding up. The k's will vary from 1 to 16